# FieldGen: From Teleoperated Pre-Manipulation Trajectories to Field-Guided Data Generation

### v1

@@ I INTRODUCTION
To overcome these challenges, we propose FieldGen, a novel field-guided data generation framework that enables scalable collection of diverse, high-quality real-world manipulation data with minimal human involvement. FieldGen is built on a key observation that manipulation tasks can be naturally decomposed into two phases with fundamentally different requirements: a pre-manipulation phase involving reaching and approaching the target object, where trajectory variation is tolerable as long as paths converge to appropriate manipulation configurations, and a fine manipulation phase requiring precise contact-rich interactions where expert demonstrations provide the most value. Our method operates as a semi-automated pipeline that leverages this decomposition by first collecting a small set of high-quality human demonstrations focused on the fine manipulation phase, particularly critical manipulation poses and contact interactions. We then construct an attraction field for the pre-manipulation phase that guides trajectories toward successful manipulation configurations identified from these human demonstrations. Using automated scripts, we generate large numbers of randomized initial observations and compute trajectories that asymptotically converge to this attraction field, producing extensive observation-action pairs for the pre-manipulation phase. By decoupling these phases, FieldGen achieves three critical benefits: high-quality supervision for fine manipulation through focused human demonstration, scalable automated generation of diverse pre-manipulation trajectories, and efficient large-scale data collection that maintains both behavioral diversity and data quality while dramatically reducing human involvement compared to traditional teleoperation approaches.
We summarize our contributions as follows: (1) Pre-manipulation field (PMF). We introduce the pre-manipulation field as a new concept, defined as an abstract representation that models and extrapolates from a small set of demonstration trajectories. PMF provides a unified framework for large-scale, semi-automated, real-robot data collection. (2) Semi-automated collection via phase decoupling. Building on PMF, we propose an efficient, semi-automated pipeline that decouples manipulation into pre-manipulation phases and fine manipulation phases, utilizing teleoperated manipulation poses to generate large-scale datasets, thereby substantially improving collection efficiency automatically. (3) Empirical validation. Experiments show that our method produces larger, higher-quality datasets in less time; models trained on our data are more robust; and operator workload is markedly reduced, enabling long-duration, stable data collection.
@@ IV-D Trajectory Diversity and Spatial Coverage
IV-E Ablation on Curve Type: Bezier vs. Cycloid
We conduct an ablation study to analyze the effect of different curve types on task success rate. In addition to the cycloid curve used in the main approach, we evaluate Bezier curves—a commonly used alternative in trajectory generation. The results are summarized in Figure 6(a). Cycloid-based trajectories achieve success rates of 75% on DP and 66.7% on RDT, corresponding to improvements of 16.7% and 8.4% over Bezier, respectively. These results highlight the clear advantage of cycloid-based trajectory generation.
To better understand this advantage, we examine the geometric properties of the two approaches. The cycloid is generated directly under the pure geometric constraint of a rolling circle without slipping. This construction provides an explicit geometric prior and yields naturally smooth curvature variations. In contrast, the Bezier curve relies on manually defined control points to shape the trajectory. Lacking inherent geometric constraints, it often exhibits sharp curvature growth over short distances, resulting in abrupt changes that increase execution difficulty for the manipulator.
@@ V CONCLUSIONS
We proposed FieldGen, a field-guided, semi-automated framework for real-world manipulation data generation. By decoupling reach and fine manipulation, and labeling reach trajectories via a cone (position) and spherical (orientation) attraction field, FieldGen converts free wrist-camera observations into high-quality ⟨obs, action⟩ pairs with minimal human effort. Real world experiments show that FieldGen consistently outperforms teleoperation under equal collection time, scales more favorably with additional data, and yields stronger generalization across start-pose, object-pose, and cross-instance settings. Ablations further indicate that cycloid-based trajectories improve success and motion smoothness over Bezier alternatives. Compared with teleoperation collection, FieldGen acquires data more efficiently while imposing substantially lower operator effort.
FieldGen thus offers a practical path to faster collection and broader coverage for large-scale real-robot datasets. Future work will extend the fields to multi-step tasks, incorporate uncertainty-aware control, and explore scene synthesis (e.g., 3DGS/NeRF) for fully automated observation generation.

### v2

@@ I INTRODUCTION
To address these challenges, we present FieldGen, a field-guided data generation framework that enables scalable collection of diverse, high-quality real-world manipulation data with minimal human effort. Our key observation is that manipulation naturally decomposes into two phases with distinct requirements: a pre-manipulation phase of reaching and approaching, where trajectory variation is acceptable as long as paths converge to valid manipulation configurations, and a fine-manipulation phase requiring precise, contact-rich control, for which expert demonstrations are most valuable. FieldGen implements a semi-automated pipeline: we first gather a small set of high-quality human demonstrations emphasizing critical poses and contact interactions. From these, we construct an attraction field that guides pre-manipulation trajectories toward the identified manipulation configurations. Automated scripts then sample randomized initial observations and compute trajectories that asymptotically converge to this field, yielding large volumes of observation–action pairs for the pre-manipulation phase. This decoupling provides three benefits: (i) high-quality supervision for fine manipulation via focused human demos, (ii) scalable automated generation of diverse pre-manipulation trajectories, and (iii) efficient, large-scale data collection that preserves both diversity and quality while greatly reducing human involvement relative to teleoperation. We further introduce FieldGen-Reward, which generates trajectories of varying quality with explicit labels, enabling models to better capture behavioral causality and improving the robustness of learned manipulation policies.
We summarize our contributions as follows: (1) Pre-manipulation field (PMF). We introduce the pre-manipulation field as a new concept, defined as an abstract representation that models and extrapolates from a small set of demonstration trajectories. PMF provides a unified framework for large-scale, semi-automated, real-robot data collection. (2) Semi-automated collection via phase decoupling. Building on PMF, we introduce a semi-automated pipeline that decouples manipulation into pre-manipulation and fine-manipulation phases: teleoperated demonstrations provide critical manipulation poses, from which we synthesize large-scale datasets for the pre-manipulation phase. The pipeline further produces trajectories of varying quality with explicit quality annotations, substantially improving automated collection efficiency while preserving diversity and fidelity. (3) Empirical validation. Experiments show that our method produces larger, higher-quality datasets in less time; models trained on our data are more robust; and operator workload is markedly reduced, enabling long-duration, stable data collection.
@@ IV-D Trajectory Diversity and Spatial Coverage
IV-E Ablation on Curve Type: B’ezier vs. Cycloid
We conduct an ablation study to analyze the effect of different curve types on task success rate. In addition to the cycloid curve used in the main approach, we evaluate B’ezier curves—a commonly used alternative in trajectory generation. The results are summarized in Figure 6(a). Cycloid-based trajectories achieve success rates of 75% on DP and 66.7% on RDT, corresponding to improvements of 16.7% and 8.4% over B’ezier, respectively. These results highlight the clear advantage of cycloid-based trajectory generation.
To better understand this advantage, we examine the geometric properties of the two approaches. The cycloid is generated directly under the pure geometric constraint of a rolling circle without slipping. This construction provides an explicit geometric prior and yields naturally smooth curvature variations. In contrast, the B’ezier curve relies on manually defined control points to shape the trajectory. Lacking inherent geometric constraints, it often exhibits sharp curvature growth over short distances, resulting in abrupt changes that increase execution difficulty for the manipulator.
@@ V CONCLUSIONS
[…]